# Synergistic Bleeding Region and Point Detection in Laparoscopic Surgical Videos

### v2

@@ 4.1 Overall Architecture
BlooDet is a dual-task collaborative detector for simultaneous bleeding region and point detection in surgical videos. As shown in Fig. 4, our framework revises the body of SAM 2 [29] by empowering edge clues to detect bleeding regions and incorporating a point branch to detect bleeding points. The whole model consists of the following processes: image encoder, mask/point memory modeling, Edge generator, mask/point decoder, and mask/point memory bank.
@@
Image Encoding. Given a set of $N$ video frames, including the current frame $I_{k}\in\mathbb{R}^{{H}\times{W}\times{3}}$ and the previous $N$-1 frames $X=\{I_{i}\}^{k-1}_{i=k-N+1}$, we first flatten frames and feed them into the image encoder inherited from SAM 2 to produce multi-scale spatial features $F\in\mathbb{R}^{{s}\times{c}\times{N}}$, where $s$ denotes the length of the feature sequence and $c$ is the feature dimension. Then, the output sequential frame features $F_{k-N},...,F_{k}$ are fed into the mask and point branches, respectively, for memory modeling.
@@
Point Branch. As shown in the bottom of Fig. 4, the point branch comprises three main parts: point memory modeling, point decoder, and point memory bank. Compared to mask memory modeling, the point memory modeling module embeds optical flow estimation to predict the displacement field between consecutive frames $[I_{k-7},...,I_{k}]$, enabling the inference of laparoscopic camera motion and viewpoint offset during surgery, thereby identifying the movement direction of current bleeding points. Further, we integrate previous mask memory features {$M^{m}_{q}$}${}_{q=k-7}^{k-1}$ from mask branch with the corresponding point features to enhance location and temporal perception as well as to narrow the search space for bleeding point coordinates. We describe this process in detail in Sec. 4.2.
@@
Afterward, the memory-enhanced point feature $F_{point}$ passes through the point decoder to predict the bleeding point. Different from the upsampling fusion in the mask decoder, we employ learnable output tokens and prompt tokens consistent with SAM 2 that interact with $F_{point}$ through self-attention and cross-attention operations [5], followed by MLP layers to predict bleeding point coordinates and confidence scores. The point memory is stored in the point memory bank for temporal modeling.
@@
Mask Branch. This branch focuses on predicting bleeding regions by mask memory modeling and coupling bleeding edge and point prompts. The top of Fig. 4 illustrates the pipeline of our mask branch. The current frame features $F_{k}$ are first fed into mask memory modeling to perform self- and cross-attention interaction [37] with mask memory features from previous frames, producing the spatial-temporal feature $F_{mask}$. After that, we introduce an edge generator that adopts multi-scale Wavelet Laplacian filters to $F_{mask}$ for edge refinement. Then, we incorporate high-resolution features from the image encoder to obtain edge maps (detailed description in Sec. 4.3). Unlike the manual intervention prompts in SAM [16, 29], we form adaptive prompt embeddings by combining the edge map $E_{m}$ from the edge generator with the point map $P_{m}$ output from point branch. Then, the prompt encoder is utilized to encode $E_{m}$ and $P_{m}$ to yield prompt features $E_{p}$ and $P_{p}$:
@@
where $\mathcal{P}$ represents prompt encoding, $\mathbf{G}$ denotes the GeLU function, Conv refers to 2$\times$2 convolution operation, and LN is layer normalization. Also, Po denotes positional encoding [34], Le stands for learned embeddings, and $\mathcal{C}[,]$ is the concatenation operation. We input prompt features along with $F_{mask}$ into the mask decoder and attain the predicted bleeding mask by upsampling and integrating with high-resolution features. Subsequently, we employ memory encoding to achieve the mask memory feature $M^{m}_{k}$ and store it in the mask memory bank. Further, the mask maps are also updated in the mask memory bank to provide spatial guidance for bleeding point detection.
Cross-branch Guidance. Our framework embraces bidirectional collaborative guidance between masks and point branches, enabling simultaneous optimization of bleeding region and bleeding point predictions. For the mask decoder, we exploit the point map produced by the point decoder as an automatic prompt input. This helps guide the decoder to focus on the target bleeding region while mitigating the interference from residual blood in the surrounding area. In point memory modeling, the predicted mask maps from previous frames can assist in optical flow estimation and in predicting the direction of the bleeding point. Besides, mask memory features are merged with point memory features to induce the point decoder to concentrate on the most likely bleeding areas while mitigating the impact of low-contrast background. With cross-branch guidance, both branches are synergistically optimized, yielding consistent bleeding region and point predictions across consecutive frames.
@@ 4.2 Point Memory Modeling in Point Branch
To detect bleeding points in consecutive frames effectively, we embed the point memory modeling module in the point branch to develop temporal clues for point features. As illustrated in Fig. 4, point memory modeling is divided into two steps: 1) combining the optical flow between consecutive frames with region maps to compensate for the relative displacement of the bleeding point caused by camera viewpoint offset, 2) interacting the average camera displacement of previous frames with mask memory features from mask branch to obtain point memory features.
For the viewpoint offset of the camera, we first utilize the frozen PWC-Net [31] for optical flow estimation. Given $N$ frames $\{I_{i}\}^{k}_{i=k-N+1}$, the optical flow $O_{i}(x,y)\in\mathbb{R}^{{H}\times{W}\times{2}}$ between two consecutive frames can be expressed as $O_{i}(x,y)=\texttt{PWC-Net}(I_{i-1},I_{i})$. Considering the instability of the optical flow in the rapidly changing bleeding region, we reverse the mask map $M_{i}$ from mask branch for each frame and combine with $O_{i}(x,y)$ to obtain the average viewpoint offset $\bar{O_{i}}(\Delta x,\Delta y)$:
@@
After that, we aggregate point memory features $M^{p}_{i}$ of previous frames in the point memory bank with $\bar{O_{i}}$ and concatenate with mask memory features $M^{m}_{i}$ to obtain the mask-guided corrected point features $\bar{F}^{ref}_{i}$. Lastly, we perform a self-attention operation on $F_{k}$ and cross-attention with $\bar{F}^{ref}_{i}$ to the memory-enhanced point feature $F_{point}$. Through the motion compensation mechanism based on optical flow estimation together with cross-branch mask guidance, we perform effective memory modeling of bleeding point features in laparoscopic scenes with camera offset.
@@ 4.3 Edge Generator in Mask Branch
Detecting bleeding regions is challenging due to the low contrast and high noise in surgical scenes. To this end, we embed a dedicated edge generator in the mask branch, which enhances the accuracy of bleeding region detection by combining multi-scale Wavelet Laplacian filters [17] with high-resolution features containing lower-level texture clues to generate edge map prompts. Concretely, we first input the spatial-temporal features $F_{mask}$ into the Gabor Wavelet Laplacian filter to enhance edge structures in the spatial domain. The Gabor Wavelet operation on the spatial position (x, y) is calculated as
$\mathcal{G}(x,y;\lambda,\theta,\psi,\sigma,\gamma)=exp{(-\frac{x^{\prime 2}+\gamma^{2}y^{\prime 2}}{2\sigma^{2}})}exp{(i(\frac{2\pi}{\lambda}x^{\prime}+\psi))}\ ,$ (5)
where $x^{\prime}=xcos\theta+ysin\theta$, $y^{\prime}=-xsin\theta+ycos\theta$, $\lambda$ denotes the wavelength, $\theta$ is the orientation angle of the Gabor kernel, $\psi$ is the phase offset, $\sigma$ is the standard deviation of Gaussian function, and $\gamma$ stands for the aspect ratio. Therefore, Laplacian filtering based on the Gabor wavelet is defined as
@@
where $\odot$ and $\ast$ denote the element-wise multiplication and convolution operation. The feature $F^{{}^{\prime}}_{mask}$ represents the output from the edge generator. As illustrated in Fig. 4, we parallel upsample $F_{mask}$ twice and separately pass through the Wavelet Laplace filters, and then interact with high-resolution features, i.e., $F_{1}$ and $F_{2}$, to further refine the edge information. Finally, the generated edge map is fed into the mask decoder as the edge prompt for bleeding region detection.
4.4 Loss Function

### v3

@@ 4.1 Overall Architecture
As shown in Fig. 5, our framework empowers edge clues to detect bleeding regions and incorporates a point branch to localize bleeding points based on SAM 2 [ravi2024sam]. The whole model consists of the following processes: image encoder, mask/point memory modeling, Edge generator, mask/point decoder, and mask/point memory bank.
@@
Image Encoding. Given a set of $N$ video frames, including the current frame $I_{k}\in\mathbb{R}^{{H}\times{W}\times{3}}$ and the previous NN-1 frames $X=\{I_{i}\}^{k-1}_{i=k-N+1}$, we first flatten frames and feed them into the image encoder inherited from SAM 2 to produce multi-scale spatial features $F\in\mathbb{R}^{{s}\times{c}\times{N}}$, where $s$ denotes the length of the feature sequence and $c$ is the feature dimension. Then, the output sequential frame features $F_{k-N},...,F_{k}$ are fed into the mask and point branches, respectively, for memory modeling.
@@
Point Branch. As shown in the bottom of Fig. 5, the point branch comprises three parts: point memory modeling, point decoder, and point memory bank. The point memory modeling module embeds optical flow estimation to predict the displacement field between consecutive frames $I_{k-7},...,I_{k}]$, enabling the inference of laparoscopic camera motion and viewpoint offset during surgery, thereby identifying the movement direction of current bleeding points. Further, we integrate previous mask memory features $M^{m}_{q}$$_{q=k-7}^{k-1}$ from mask branch with the corresponding point features to enhance location and temporal perception as well as to narrow the search space for bleeding point coordinates. We describe this process in detail in Sec. 4.2.
@@
Afterward, the memory-enhanced point feature $F_{point}$ passes through the point decoder to predict the bleeding point. Different from the upsampling fusion in the mask decoder, we employ learnable output tokens and prompt tokens that interact with $F_{point}$ via self-attention and cross-attention [carion2020end], followed by MLP layers to predict point coordinates and confidence scores. The point memory is stored in the point memory bank for temporal modeling.
@@
Mask Branch. The top of Fig. 5 illustrates the pipeline of the mask branch, which produces bleeding regions by mask memory modeling and coupling edge and point prompts. The current frame features $F_{k}$ are first fed into mask memory modeling to perform self- and cross-attention interaction [vaswani2017attention] with mask memory features from previous frames, producing the spatial-temporal feature $F_{mask}$. After that, we introduce an edge generator that adopts multi-scale wavelet Laplacian filters to $F_{mask}$ for edge refinement. Then, we incorporate high-resolution features from the image encoder to obtain edge maps (detailed description in Sec. 4.3). Unlike the manual intervention prompts [kirillov2023segment, ravi2024sam], we form adaptive prompt embeddings by combining the edge map $E_{m}$ from the edge generator with the point map $P_{m}$ from point branch. Then, the prompt encoder is utilized to yield prompt features $E_{p}$ and $P_{p}$:
@@
where $\mathcal{P}$ represents prompt encoding, $\mathbf{G}$ denotes the GeLU function, Conv refers to $\times2$ convolutions, and LN is layer normalization. Also, Po denotes positional encoding [tancik2020fourier], Le stands for learned embeddings, and $\mathcal{C}[,]$ is the concatenation operation. We input prompt features along with $F_{mask}$ into the mask decoder and attain the bleeding mask by upsampling and integrating with high-resolution features. Then, we employ memory encoding to achieve mask memory feature $M^{m}_{k}$ and store it in the mask memory bank. The mask maps are also updated in the memory bank to provide spatial guidance for bleeding point detection.
Cross-branch Guidance. Our framework embraces bidirectional collaborative guidance between masks and point branches, enabling simultaneous optimization of bleeding region and bleeding point predictions. In mask decoder, we exploit the point map produced by the point decoder as an automatic prompt input. This helps guide the decoder to focus on the target bleeding region while mitigating the interference from residual blood in the surrounding area. In point memory modeling, the predicted mask maps from previous frames can assist in optical flow estimation and in predicting the direction of the bleeding point. Besides, mask memory features are merged with point memory features to induce the point decoder to concentrate on the most likely bleeding areas while mitigating the impact of low-contrast background. The alternating optimization strategy for cross-branch guidance is found in Sec. 4.4.
@@ 4.2 Point Memory Modeling in Point Branch
To detect bleeding points in consecutive frames effectively, we embed the point memory modeling module in the point branch to develop temporal clues for point features. As illustrated in Fig. 5, point memory modeling is divided into two steps: 1) combining the optical flow between consecutive frames with region maps to compensate for the relative displacement of the bleeding point caused by camera viewpoint offset, as shown in Fig. 6; 2) interacting the average camera displacement of previous frames with mask memory features from mask branch to obtain point memory features.
Figure 6: Visualization of optical flow–based camera compensation for the relative motion of the bleeding point. Despite the flow in operative areas being complex, the overall flow direction toward the upper left (blue-purple), consistent with the point motion.
For the viewpoint offset of the camera, we first utilize the frozen PWC-Net [sun2018pwc] for optical flow estimation. Given $N$ frames $\{I_{i}\}^{k}_{i=k-N+1}$, the optical flow $O_{i}(x,y)\in\mathbb{R}^{{H}\times{W}\times{2}}$ between two consecutive frames can be expressed as $O_{i}(x,y)=\texttt{PWC-Net}(I_{i-1},I_{i})$. Considering the instability of the optical flow in the rapidly changing bleeding region, we reverse the mask map $M_{i}$ from mask branch for each frame and combine with $O_{i}(x,y)$ to obtain the average viewpoint offset $\bar{O_{i}}(\Delta x,\Delta y)$:
@@
After that, we aggregate point memory features $M^{p}_{i}$ of previous frames in the point memory bank with $\bar{O_{i}}$ and concatenate with mask memory features $M^{m}_{i}$ to obtain the mask-guided corrected point features $\bar{F}^{ref}_{i}$. Lastly, we perform self-attention on $F_{k}$ and cross-attention with $\bar{F}^{ref}_{i}$ to the memory-enhanced point feature $F_{point}$. Through the optical flow estimation as well as mask guidance, we model the effective memory of point traits with camera offset.
@@ 4.3 Edge Generator in Mask Branch
To address low contrast and high noise in surgical fields, we embed an edge generator in the mask branch that generates edge map prompts by combining multi-scale wavelet Laplace filters [lee1996image] with high-resolution features containing low-level texture cues. Concretely, we first input the spatial-temporal features $F_{mask}$ into the Gabor wavelet Laplacian filter to enhance edge structures in the spatial domain. The Gabor wavelet operation on position (x, y) is calculated as
$\small\mathcal{G}(x,y;\lambda,\theta,\psi,\sigma,\gamma)=exp{(-\frac{x^{\prime 2}+\gamma^{2}y^{\prime 2}}{2\sigma^{2}})}exp{(i(\frac{2\pi}{\lambda}x^{\prime}+\psi))}\ ,$ (6)
where $x^{\prime}=xcos\theta+ysin\theta$, $y^{\prime}=-xsin\theta+ycos\theta$, $\lambda$ denotes the wavelength, $\theta$ is the orientation angle of the Gabor kernel, $\psi$ is the phase offset, $\sigma$ is the standard deviation of Gaussian function, and $\gamma$ stands for the aspect ratio. Thus, Laplacian filtering based on the Gabor wavelet is defined as
@@
where $\odot$ and $\ast$ denote the element-wise multiplication and convolution operation. The feature $F^{{}^{\prime}}_{mask}$ represents the output from the edge generator. As illustrated in Fig. 5, we parallel upsample $F_{mask}$ twice and separately pass through the wavelet Laplacian filter, and then interact with high-resolution features ($F_{1}$ and $F_{2}$) to refine bleeding edges. Finally, the generated edge map is fed into the mask decoder as prompts for bleeding region detection.
Table 2: Overall comparison with task-related methods on SurgBlood test set and HemoSet [miao2024hemoset] dataset. Scores are reported as mean $\pm$ standard deviation. $\dagger$ and * denote region- and point-level models with an additional point/region prediction head for dual-task learning.
Methods Volumes Bleeding Region Detection Bleeding Point Detection Params $\downarrow$
SurgBlood HemoSet [miao2024hemoset] SurgBlood
IoU $\uparrow$ Dice $\uparrow$ IoU $\uparrow$ Dice $\uparrow$ PCK-2% $\uparrow$ PCK-5% $\uparrow$ PCK-10% $\uparrow$
Swin-UNet [swinunet] ECCV’22 $34.83\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.15}}$ $51.67\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.20}}$ $29.86\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.12}}$ $45.99\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.17}}$ $65\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.12}} 11.$$1.70\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.18}}$ 37.23 ± $3\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.22}} 27$.2M
SAM [kirillov2023segment] ICCV’23 $37.94\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.17}}$ $55.01\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.23}}$ $38.81\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.18}}$ $55.91\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.21}}$ $4.61\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.14}}$ $25.86\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.19}}$ $60.79\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.21}}$ 93.9M
MemSAM [deng2024memsam] CVPR’24 $52.84\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.16}}$ $69.14\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.22}}$ $46.89\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.17}}$ $63.84\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.24}}$ $5.27\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.13}}$ $31.80\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.24}}$ $64.91\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.18}}$ 133.5M
STDDNet [chen2025stddnet] ICCV’25 $47.49\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.18}}$ $64.39\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.19}}$ $43.98\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.16}}$ $61.10\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.26}}$ $1.48\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.14}}$ $13.67\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.17}}$ $45.80\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.20}}$ 38.2M
SAM 2 [ravi2024sam] ICLR’25 $50.93\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.12}}$ $67.49\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.14}}$ $56.53\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.19}}$ $72.02\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.24}}$ $12.35\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.17}}$ $41.68\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.21}}$ $71.99\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.26}}$ 81.0M
HRNet [sun2019deep] CVPR’19 $48.90\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.12}}$ $65.68\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.17}}$ $48.07\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.18}}$ $64.93\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.26}}$ $1.98\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.15}}$ $15.98\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.18}}$ $47.28\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.23}}$ 63.6M
SimCC [li2022simcc] ECCV’22 $51.02\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.17}}$ $67.57\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.20}}$ $51.62\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.19}}$ $68.09\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.23}}$ $3.62\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.13}}$ $17.30\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.20}}$ $48.76\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.16}}$ 66.3M
GTPT [wang2024gtpt] ECCV’24 $41.10\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.14}}$ $58.26\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.18}}$ $50.27\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.20}}$ $66.91\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.22}}$ $3.29\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.11}}$ $15.98\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.25}}$ $38.55\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.18}}$ 16.7M
D-CeLR [dai2024cephalometric] ECCV’24 $51.30\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.19}}$ $67.82\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.19}}$ $50.85\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.21}}$ $67.42\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.25}}$ $2.97\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.12}}$ $24.22\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.24}}$ $63.92\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.17}}$ 53.4M
CalibratedSL [feng2025uncertainty] TMI’25 $36.96\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.16}}$ $53.97\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.16}}$ $44.59\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.16}}$ $61.68\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.19}}$ $3.46\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.14}}$ $15.82\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.22}}$ $39.70\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.13}}$ 6.6M
PAINet [das2023multi] MICCAI’23 $44.14\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.14}}$ $61.24\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.15}}$ $53.78\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.17}}$ $69.87\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.21}}$ $2.47\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.19}}$ $15.48\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.26}}$ $48.43\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.20}}$ 13.6M
PitSurgRT [mao2024pitsurgrt] IJCARS’24 $30.48\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.15}}$ $46.72\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.17}}$ $27.11\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.18}}$ $42.65\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.24}}$ $2.47\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.18}}$ $13.84\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.23}}$ $41.68\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.18}}$ 67.3M
ConsisTNet [mao2025consistnet] IJCARS’25 $40.43\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.16}}$ $57.59\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.18}}$ $49.76\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.22}}$ $66.45\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.26}}$ $7.09\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.15}}$ $32.83\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.19}}$ $68.15\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.21}}$ 143.5M
[…]
4.4 Objective Function
Synergizing two tasks within a unified model to solve coupled objectives is nontrivial, as updating one set of parameters immediately changes the optimal solution for the other. To this end, we adopt an alternating optimization scheme: fixing one branch during iteration while iteratively updating the other. At iteration $t$, we perform two coordinated steps:
1) Update region detection given the current point branch. We first fix $\bm{\vartheta}^{(t)}$ and update $\bm{\theta}$ by minimizing the region detection loss, which is conditioned on (and regularized by) the current point branch:
$\bm{\theta}^{(t+1)}=\bm{\theta}^{(t)}-\eta_{\theta}\,\nabla_{\bm{\theta}}\,\mathcal{L}_{\mathtt{m}}\Big(\bm{\theta}^{(t)}\big(\bm{\vartheta}^{(t)}\big)\Big).$ (10)
2) Update point localization given the updated detector. We then fix $\bm{\theta}^{(t+1)}$ and update $\bm{\vartheta}$ by minimizing the localization loss under the latest detector context:
$\bm{\vartheta}^{(t+1)}=\bm{\vartheta}^{(t)}-\eta_{\vartheta}\,\nabla_{\bm{\vartheta}}\,\mathcal{L}_{\mathtt{p}}\Big(\bm{\vartheta}^{(t)}\big(\bm{\theta}^{(t+1)}\big)\Big).$ (11)
This alternating strategy enforces mutual adaptation: the detector branch $\bm{\theta}$ is optimized under fine-grained geometric constraints from the current localization branch $\bm{\vartheta}^{(t)}$, while the localization branch $\bm{\vartheta}$ is optimized within the updated spatial context provided by $\bm{\theta}^{(t+1)}$. Repeating Eq. (10) and Eq. (11) drives both sets of parameters toward a joint fixed point that is consistent across region- and point-level predictions, thus approximating the coupled optimum in Eq. (1).